# TextlessRAG: End-to-End Visual Document RAG by Speech Without Text

### v1

@@ 2 Method
As shown in Figure 2, TextLessRAG is designed to be both concise and efficient focusing specifically on the retrieval and QA generation components. On Retrieval side, we employ ColQwen-Omni [^1] https://huggingface.co/vidore/colqwen-omni-v0.1 as a retrieve encoder $\mathrm{Enc}(\cdot)$ to encode both document image pages from document knowedge base $\mathcal{I}=\{P_{1},P_{2}...P_{n}\}$and spoken query $q$ to embeddings as $E=\{e_{1},e_{2}...e_{n}\}$ and $e_{q}$. Then compute the MaxSim Score like ColBert [^2] https://github.com/stanford-futuredata/ColBERT to rank the top-k candidate pages $T_{k}$:
@@
On Generation side, we employ Qwen2.5-Omni [^3] https://huggingface.co/Qwen/Qwen2.5-Omni-7B as a Generator $Gen(\cdot)$, which directly takes the spoken query $q$ together with the top-k re-ranked document images $\mathcal{T}_{k}^{\prime}$ as input, and ultimately produces a spoken answer $\mathcal{A}ns$:
@@
We lay out each top-k page $P_{t}$ into content blocks of chart, table, text and natural image sets by DocLayout-Yolo [^4] https://github.com/opendatalab/DocLayout-YOLO:
@@ 3 Data Engine
Specifically, (1) we collect a diverse set of documents and reports across multiple domains and formats, including PDFs and PPTs (see the domain word cloud in Figure 3, left). (2) Each page is segmented into fine-grained units—tables, charts, images, and text blocks—using DocLayout-YOLO. (3) These segments are processed by commercial vision–language agents via batch APIs to generate candidate QA pairs. (4) The QA pairs are refined through rule-based filtering and professional human annotation. (5) Finally, the text QA pairs are converted into speech using Doubao’s TTS API, leveraging over 200 professional voice types. [^5] https://www.volcengine.com/docs/6561/1257544.
@@ 4.1 Retrival Result
Table 2 reports Retrieval results, comparing three text-only methods—BM25[16], E5[17], and NV-Embed-v2[18]— against five visual models(with ground truth text input as query) — CLIP[19], DSE[20], VisRAG-Ret[8], VDocRAG[9], and ViDoRAG[10]—as baselines. For text-based retrieval, OCR text was extracted using Tesseract [^6] https://github.com/tesseract-ocr/tesseract. Retrieval experiments were then conducted independently within each retrieval pool across seven datasets.

### v2

# Index Terms—  Visual Document Understanding, Audio Visual Question Answering, Multimodal RAG

@@ 2 Method
As shown in Figure 2, TextLessRAG is designed to be both concise and efficient focusing specifically on the retrieval and QA generation components. On Retrieval side, we employ ColQwen-Omnihttps://huggingface.co/vidore/colqwen-omni-v0.1 as a retrieve encoder $\mathrm{Enc}(\cdot)$ to encode both document image pages from document knowedge base $\mathcal{I}=\{P_{1},P_{2}...P_{n}\}$and spoken query $q$ to embeddings as $E=\{e_{1},e_{2}...e_{n}\}$ and $e_{q}$. Then compute the MaxSim Score like ColBerthttps://github.com/stanford-futuredata/ColBERT to rank the top-k candidate pages $T_{k}$:
@@
On Generation side, we employ Qwen2.5-Omnihttps://huggingface.co/Qwen/Qwen2.5-Omni-7B as a Generator $Gen(\cdot)$, which directly takes the spoken query $q$ together with the top-k re-ranked document images $\mathcal{T}_{k}^{\prime}$ as input, and ultimately produces a spoken answer $\mathcal{A}ns$:
@@
We lay out each top-k page $P_{t}$ into content blocks of chart, table, text and natural image sets by DocLayout-Yolohttps://github.com/opendatalab/DocLayout-YOLO:
@@ 3 Data Engine
Specifically, (1) we collect a diverse set of documents and reports across multiple domains and formats, including PDFs and PPTs (see the domain word cloud in Figure 3, left). (2) Each page is segmented into fine-grained units—tables, charts, images, and text blocks—using DocLayout-YOLO. (3) These segments are processed by commercial vision–language agents via batch APIs to generate candidate QA pairs. (4) The QA pairs are refined through rule-based filtering and professional human annotation. (5) Finally, the text QA pairs are converted into speech using Doubao’s TTS API, leveraging over 200 professional voice types.https://www.volcengine.com/docs/6561/1257544.
@@ 4.1 Retrival Result
Table 2 reports Retrieval results, comparing three text-only methods—BM25[16], E5[17], and NV-Embed-v2[18]— against five visual models(with ground truth text input as query) — CLIP[19], DSE[20], VisRAG-Ret[8], VDocRAG[9], and ViDoRAG[10]—as baselines. For text-based retrieval, OCR text was extracted using Tesseracthttps://github.com/tesseract-ocr/tesseract. Retrieval experiments were then conducted independently within each retrieval pool across seven datasets.